OpenStudy (anonymous):

@jim_thompson5910 xcan you help me please on finding the points

OpenStudy (anonymous):

Graph the circle (x - 2) 2 + (y + 1) 2 = 48.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (campbell_st):

the general form of a circle is \[(x -h)^2 + (y -k)^2 = r^2\] so you have a circle with a centre od (2, -1) and the radius is \[\sqrt{48} ... or.... 4\sqrt{3}\] providing its not a typo and should be 49

OpenStudy (anonymous):

@campbell_st can you help me?

OpenStudy (campbell_st):

well is it a typo or no.... is it 48 or 49..?

OpenStudy (anonymous):

its 49

OpenStudy (anonymous):

yay jims here :)

jimthompson5910 (jim_thompson5910):

hey, sry was helping someone else out, but it looks like campbell_st has got it

OpenStudy (anonymous):

i still dont get it

jimthompson5910 (jim_thompson5910):

but yes, this is a circle with a center of (2,-1) and has a radius of 7 (since the square root of 49 is 7)

OpenStudy (campbell_st):

ok... so the radius of you circle is 7 and the centre is (2, -1) so plot the centre of the circel on the number plane

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

so you start at the point (2,-1) then you draw a circle around that point that has a radius of 7

OpenStudy (anonymous):

so i go 7 up, dwn,left,and right?

jimthompson5910 (jim_thompson5910):

yep you start at (2,-1) and you go 7 units in every direction

jimthompson5910 (jim_thompson5910):

to plot points that lie on the circle

OpenStudy (anonymous):

awesome thank you both :D

OpenStudy (campbell_st):

then you will need to plot 4 points (2, 6) (2, -8) (9, 6) and (-5, 6) these points are on the circumference... and with the centre will divide the circle into quadrants... just join the circumference points with arcs.